A given memory system has a page size of 1 KB. If a process references memory location 9210, calculate the page number and the offset within the page. (Show all work)
Answer #1

Each page is of size 1KB = 1024 bytes/locations .

Given the address /location = 9210

Divide the address by 1024 to get the no of pages/page number = 9210/1024 = 8.994140625 = here integer part is page number and decimal part is the offset in that page itself

So page number = 8

Offset = multiply by 1024 the decimal part = 1024×0.994140625= 1018 = offset in page 8

We can verify using formula :

Address location = page number*1024 +(offset)

= 1024×8 + 1018 = 8092 + 1018

= 9210 = hence proved